OpenStudy (anonymous):

An elevator is going down with a constant acceleration. A coin is dropped from a point 1.8m above the floor takes 1 second to reach the floor The magnitude of the acceleration of the elevator is ? (g=10m/s^)

OpenStudy (apoorvk):

so, the net acceleration of the coin g minus the acc. of the lift. use this funda.

OpenStudy (anonymous):

Relative Acceleration = g-a

OpenStudy (anonymous):

Exactly. Now that you are in the lift. It's a normal question again. s = 1.8 t = 1s a = g - x ( x is acceleration of lift) s = 1/2at^2.

OpenStudy (anonymous):

ok the answer is 6.4m/s^2 is it @siddhantsharan

OpenStudy (anonymous):

Yeah. I think so.

OpenStudy (anonymous):

Thxxx

OpenStudy (aravindg):

exactlly the question talks of a non-inertial system, only thing u have to do is to make it to a inertial system by using a pseudo force ,the rest of ur work is simply applying newttons eqns !!